A ladder 5 meter long leans against the wall of an apartment house forming an angle of 50 degrees, 32 minutes with the ground. how high up the wall does it reach?
Answer #1
sin theta = H / L = H / 5

H = 5 * sin theta = 5 * sin 50 ° 32 ´

H = 5 * 0.77199

H = 3.85995 m approx. H = 3.86 m
answered by: LauRen Burgess